Is log(x!) differentiable ? If yes then dlog(x!)/dx = ?

Grade:12th pass

1 Answers

Nandana
7 years ago
yes , it is …
    firstly , d/dx (x!) = Gamma (x+1) ;
       take   x!= u , then
         d/du (log(u)) = 1/u
        so , d/dx (log x!) = 1/x!* d/dx(x!)
                                   = Gamma (x+1) /x!